32.00 dollars total bill. Jeans cost 10.00 dollars more than shirt. What did jeans cost?
Answer by benazir.sj@gmail.com(109)

cost of shirt =x
cost of jeans =x+10
total cost =$32
x+x+10 =32
2x+10 =32
2x=32-10
2x=22
x=11
cost of shirt =$11
cost of jeans =x+10
=$11+$10
=$21